In order that work may be done,
1
Expert's answer
2020-12-09T09:48:31-0500

Answer

Work done is given by

"dW=\int F.ds"

If force is constant then work done given by

"W=Fdcos\theta"


In order to work done zero

"\theta" should be 90° or force and displacement should be perpendicular

So

"W=Fdcos90°\\W=0J"